OpenStudy (help!!!!):

10. A student receives his grade report from a local community college, but the GPA is smudged. He took the following classes: a 2 hour credit art, a 3 hour credit history, a 4 hour credit science course, a 3 hour credit mathematics course, and a 1 hour science lab. He received a “B” in the art class, an “A” in the history class, a “C” in the science class, a “B” in the mathematics class, and an “A” in the science lab. What was his GPA if the letter grades are based on a 4 point scale? (A=4, B=3, C=2, D=1, F=0) A. 2.7 B. 2.8 C. 3.0 D. 3.1 E. 3.2

OpenStudy (help!!!!):

steps would be appreciated

OpenStudy (anonymous):

Okay, so you assign the numbers to the letter grades and multiply times the number of credits for each class. In the first case, you have a "B" in a three unit class. B= 3.0, so you mutiply that times the credits 3*3 =9. Do that for the remaining classes, and add up the point totals. Then divide by the number of hours to get the GPA.

OpenStudy (anonymous):

Ooops the first one was only two credits, so it should be 6..

OpenStudy (anonymous):

(2*3)+(3*4)+(4*2)+(3*3)+(4*1)= 6+12+8+9+4= 39 13 units total, so the GPA is 39/13= 3.0

OpenStudy (help!!!!):

thanks